# Quantum Sieving for Code-Based Cryptanalysis and Its Limitations for ISD

### v1

@@ Abstract
Sieving using near-neighbor search techniques is a well-known method in lattice-based cryptanalysis, yielding the current best runtime for the shortest vector problem in both the classical [BDGL16] and quantum [BCSS23] setting. Recently, sieving has also become an important tool in code-based cryptanalysis. Specifically, using a sieving subroutine, [GJN23, DEEK24] presented a variant of the information-set decoding (ISD) framework, which is commonly used for attacking cryptographically relevant instances of the decoding problem. The resulting sieving-based ISD framework yields complexities close to the best-performing classical algorithms for the decoding problem such as [BJMM12, BM18]. It is therefore natural to ask how well quantum versions perform. In this work, we introduce the first quantum algorithms for code sieving by designing quantum variants of the aforementioned sieving subroutine. In particular, using quantum-walk techniques, we provide a speed-up over the best known classical algorithm from [DEEK24] and over a variant using Grover’s algorithm [Gro96]. Our quantum-walk algorithm exploits the structure of the underlying search problem by adding a layer of locality-sensitive filtering, inspired by the quantum-walk algorithm for lattice sieving from [CL21]. We complement our asymptotic analysis of the quantum algorithms with numerical results, and observe that our quantum speed-ups for code sieving behave similarly as those observed in lattice sieving. In addition, we show that a natural quantum analog of the sieving-based ISD framework does not provide any speed-up over the first presented quantum ISD algorithm [Ber10]. Our analysis highlights that the framework should be adapted in order to outperform the state-of-the-art of quantum ISD algorithms [KT17, Kir18].
@@ 1.1 Main Contributions
We present several quantum algorithms for FindSolutions and analyze their asymptotic time and memory complexities. Our first quantum algorithm is a straightforward application of Grover, and serves as a baseline for comparison. Our other quantum algorithms are a quantum-walk algorithm and a variant thereof, where we apply the sparsification technique from [CL21]. Both quantum-walk algorithms make use of locality-sensitive filtering (LSF) to obtain a speed-up over Grover in the search for ‘solutions’. The use of LSF inside the quantum walk was first suggested by [CL21] in the context of lattice sieving, and one of our contributions is to show how to apply this layer of LSF in the context of codes (see Section 4.4). Specifically, we introduce a Hamming-metric variant of the ‘residual vectors’ used in [CL21], which behave somewhat differently than their Euclidean-metric analogs, but still allow us to apply LSF.
Besides providing an asymptotic analysis of the time and memory complexities of our quantum algorithms, we evaluate their performance through numerical optimization. Our numerical results illustrate the obtained quantum speed-up over the classical algorithm from [DEEK24], as well as a comparison between our different quantum algorithms. In addition, we show that the quantum speed-ups we obtain align with the speed-ups observed in the state-of-the-art of lattice cryptanalysis (see Table 1). The Python code used for the numerical results of this work is available at https://github.com/lynnengelberts/quantum-sieving-for-codes-public.
@@ 2.4 Quantum Computation
We say that a quantum algorithm has time complexity $T$ if the circuit describing it uses at most $T$ elementary quantum gates. When $T$ is polynomial in the number of bits needed to write down the input, we say the algorithm is efficient.
@@
For $N\in\mathbb{N}$, let $f\colon[N]\to\{0,1\}$ be a function and define $t:=|\{i\in[N]:f(i)=1\}|$. There exists a quantum algorithm, called Grover’s algorithm, that returns $i$ such that $f(i)=1$, if such an $i$ exists, using $O(\sqrt{N/\max\{1,t\}})$ queries to $f$ and $\tilde{O}(\sqrt{N/\max\{1,t\}})$ elementary gates with probability at least $2/3$. It uses $\max(\log_{2}(x_{i}))$ quantum memory and requires efficient QRAM access to the register $x\in\{0,1\}^{N}$ with $x_{i}=f(i)$. In particular, if the register is classically stored, it requires QRACM of size $N$, and if it is quantumly stored, QRAQM of size $N$.
@@
For $N\in\mathbb{N}$, let $f\colon[N]\to\{0,1\}$ be a function. Suppose there is an efficient quantum circuit $U_{f}$ that maps |$\ket{i}\to(-1)^{f(i)}\ket{i}$. Suppose $\mathcal{A}$ is an algorithm that can be implemented as a reversible quantum circuit and that returns $i\in[N]$ such that $f(i)=1$ in time $T$ with success probability $p>0$. Then there exists a quantum algorithm, called amplitude amplification, that returns $i\in[N]$ such that $f(i)=1$ with probability at least $\max(1-p,p)$ in time $O(1/\sqrt{p})\cdot T$. If algorithm $\mathcal{A}$ uses quantum access to a classical register (resp. quantum), then running amplitude amplification requires QRACM (resp. QRAQM) operators.
@@ 3.1 Framework for Code Sieving
1 Sample a tower of codes $\{\mathbb{F}_{2}^{n}=\mathcal{C}_{0},\dots,\mathcal{C}_{n-k}=\mathcal{C}\}$, with dimension decrements of 1.
2 Sample $N$ vectors independently and uniformly at random from $\mathcal{S}_{w}^{n}$ and add them to a list $\mathcal{L}_{0}$.
@@
4       Invoke $\mathcal{A}_{\textsc{NNS}}$ on $\mathcal{L}_{i-1}$ to obtain $\mathcal{L}^{\prime}_{i}:=\{(\mathbf{x},\mathbf{y})\colon\mathbf{x},\mathbf{y}\in\mathcal{L}_{i-1},|\mathbf{x}+\mathbf{y}|=w\}$.
@@
7                   Add $\mathbf{x}+\mathbf{y}$ to $\mathcal{L}_{i}$.
@@
9      Discard some elements if $|\mathcal{L}_{i}|>N$.
@@
By Lemma 3.1, if $N\geq{n\choose w}\big{/}\big{(}{w\choose w/2}{n-w\choose w/2}\big{)}$, then $\textsc{NNS}(n,w,N)$ has at least $N$ solutions in expectation. However, note that, in each iteration $i$ of Algorithm 1, on average only a quarter of the NNS solutions $(\mathbf{x},\mathbf{y})$ found by $\mathcal{A}_{\textsc{NNS}}$ will be added to the new list $\mathcal{L}_{i}$. Indeed, half of them is discarded as each pair is counted twice: $(\mathbf{x},\mathbf{y})$ and $(\mathbf{y},\mathbf{x})$. Secondly, the algorithm discards those $\mathbf{x}+\mathbf{y}$ that do not belong to $\mathcal{C}_{i}$. Since two subsequent codes $\mathcal{C}_{i-1},\mathcal{C}_{i}$ differ by one dimension, it follows that in expectation another half of the NNS solutions are discarded. Therefore, [DEEK24] propose to take $N\geq 4\cdot{n\choose w}\big{/}\big{(}{w\choose w/2}{n-w\choose w/2}\big{)}$ to maintain a list size of $N$ through all iterations in the sieving part.
@@ 3.2 NNS using Locality-Sensitive Filtering (LSF)
The underlying idea of locality-sensitive filtering (in $\mathbb{F}_{2}^{n}$) is the following: vectors $\mathbf{x},\mathbf{y}$ satisfying $|\mathbf{x}+\mathbf{y}|=w$, also known as near neighbors, can be found more efficiently if we restrict our search to local regions of $\mathbb{F}_{2}^{n}$. The algorithm proceeds as follows. It starts with covering the space $\mathbb{F}_{2}^{n}$ with (potentially overlapping) regions, each region corresponding to a certain center $\mathbf{c}\in\mathcal{C}_{f}$, for some subset $\mathcal{C}_{f}\subseteq\mathbb{F}_{2}^{n}$. Each vector $\mathbf{x}$ from the input list $\mathcal{L}$ is then filtered according to the centers, namely, it is inserted into a bucket corresponding to a center $\mathbf{c}$ if and only if $|\mathbf{x}\land\mathbf{c}|=\alpha$ [^4] The analog in the lattice setting is that a unit vector is added to a bucket corresponding to some unit vector $\mathbf{c}\in\mathbb{R}^{n}$ if it has (absolute) inner product at least $\alpha$ with $\mathbf{c}$. This is what we refer to as the bucketing phase. Each bucket thus contains all the elements from $\mathcal{L}$ that belong to the same region. The algorithm then searches for near neighbors within each bucket, which we refer to as the checking phase. As the checking phase significantly affects the overall cost of the algorithm, our primary goal will be to improve its runtime by quantizing this part of the algorithm.
@@
The resulting NNS algorithm is presented in Algorithm 2. It uses two subroutines: FindValidCenters and FindSolutions. For a given $\mathbf{x}\in\mathcal{L}$ and bucketing parameter $\alpha$, the FindValidCenters subroutine returns a set $\mathcal{VC}_{\alpha}(\mathbf{x})=\{\mathbf{c}\in\mathcal{C}_{f}\ |\ |\mathbf{x}\land\mathbf{c}|=\alpha\}$. For a suitable choice of $\mathcal{C}_{f}$ (using so-called random product codes), the time complexity of this subroutine has been shown by [DEEK24] to be $|\mathcal{VC}_{\alpha}(\mathbf{x})|+2^{o(n)}$, as explained in the next section.
@@
3       $\mathcal{B}_{\alpha}(\mathbf{c})=\emptyset$.
@@
7             Add $\mathbf{x}$ to $\mathcal{B}_{\alpha}(\mathbf{c})$.
@@
9 $\mathcal{L}^{\prime}=\emptyset$.
@@
11       Add $\textnormal{{FindSolutions}}(\mathcal{B}_{\alpha}(\mathbf{c}))$ to $\mathcal{L}^{\prime}$.
@@
Depending on the size of $|\mathcal{C}_{f}|$, Algorithm 2 has to be repeated a certain number of times to find all solutions. The size of $|\mathcal{C}_{f}|$ also affects other factors of the time and memory complexity of Algorithm 2, and should thus be chosen carefully. The best time and memory complexity [DEEK24, Cor. 4.2] is obtained by choosing $|\mathcal{C}_{f}|$ such that $\mathbb{E}[|\mathcal{VC}_{\alpha}(\mathbf{x})|]=2^{o(n)}$, ensuring that the expected runtime of FindValidCenters is $2^{o(n)}$. (An algorithm with similar time and memory complexities as [DEEK24, Cor. 4.2] was independently obtained by Carrier [Car20, Cor. 8.2.6].)
@@ 3.3 Random Product Codes and an Efficient Algorithm for FindValidCenters
The set of centers is then sampled uniformly at random from $R_{n,v,t,\kappa}$ to guarantee: {outline} \1[(1)] Efficient decodability: in some reasonable parameter regimes, one can compute $\mathcal{VC}_{\alpha}(\mathbf{x}):=\{\mathbf{c}\in\mathcal{C}\colon|\mathbf{x}\land\mathbf{c}|=\alpha\}$ in time asymptotically equal to its size $|\mathcal{VC}_{\alpha}(\mathbf{x})|$. \1[(2)] Random behavior: for $t$ not too large, a sample $\mathcal{C}$ behaves like a fully random code in the sense that the success probability that $\mathcal{C}\sim R_{n,v,t,\kappa}$ ‘captures’ a pair $(\mathbf{x},\mathbf{y})\in\mathcal{L}^{2}$ is the same (up to factors subexponential in $n$) as for a fully random code in $S_{v}^{n}$. [^5] This ‘randomness’ property puts a constraint on $t$. More precisely, as described in [DEEK24] (and in [BDGL16] for $\mathbb{R}^{n}$), we would like $t$ to be small enough to guarantee that we can approximate a cap/wedge in $\mathbb{F}_{2}^{n}$ by the Cartesian product of $t$ caps/wedges in $\mathbb{F}_{2}^{n/t}$. By Lemma 4.6 and Lemma 4.7 in [DEEK24], these approximations are satisfied up to a subexponential factor in $n$ if $t=o(n/\log(n))$.
@@ 4 Quantum Algorithms for NNS and Code Sieving
We recall the context in which FindSolutions is used. That is, consider the successful completion of the bucketing phase in Algorithm 2 on an input list $\mathcal{L}\subseteq\mathcal{S}_{w}^{n}$: for some bucketing parameter $\alpha$, we have sampled a set $\mathcal{C}_{f}\subseteq\mathcal{S}_{v}^{n}$ of bucket centers and an associated data structure containing for each $\mathbf{c}\in\mathcal{C}_{f}$ the bucket
$\mathcal{B}_{\alpha}(\mathbf{c}):=\{\mathbf{x}\in\mathcal{L}\colon|\mathbf{x}\land\mathbf{c}|=\alpha\}\subseteq\mathrm{Region}_{\alpha}(\mathbf{c}):=\{\mathbf{x}\in\mathcal{S}_{w}^{n}:|\mathbf{x}\wedge\mathbf{c}|=\alpha\}.$ (1)
Then the goal of FindSolutions is to find all $\mathbf{x},\mathbf{y}\in\mathcal{B}_{\alpha}(\mathbf{c})$ satisfying $|\mathbf{x}+\mathbf{y}|=w$.
@@
Note that this problem is a variant of the near-neighbor search problem (Problem 3.2) defined in Section 3. An important difference here is that the input list $\mathcal{B}_{\alpha}(\mathbf{c})$ is not uniformly random on $\mathcal{S}_{w}^{n}$ but on the set $\{\mathbf{x}\in\mathcal{S}_{w}^{n}:|\mathbf{x}\wedge\mathbf{c}|=\alpha\}$. Consequently, the probability (denoted $p$ below) that a uniformly random pair from the input list forms a solution pair does not only depend on $n,w$, but also on the parameters $v,\alpha$, as further shown in the next section.

### v2

@@ Abstract
Sieving using near-neighbor search techniques is a well-known method in lattice-based cryptanalysis, yielding the current best runtime for the shortest vector problem in both the classical [BDGL16] and quantum [BCSS23] setting. Recently, sieving has also become an important tool in code-based cryptanalysis. Specifically, using a sieving subroutine, [GJN23, DEEK24] presented a variant of the information-set decoding (ISD) framework, which is commonly used for attacking cryptographically relevant instances of the decoding problem. The resulting sieving-based ISD framework yields complexities close to the best-performing classical algorithms for the decoding problem such as [BJMM12, BM18]. It is therefore natural to ask how well quantum versions perform. In this work, we introduce the first quantum algorithms for code sieving by designing quantum variants of the aforementioned sieving subroutine. In particular, using quantum-walk techniques, we provide a speed-up over the best known classical algorithm from [DEEK24] and over a variant using Grover’s algorithm. Our quantum-walk algorithm exploits the structure of the underlying search problem by adding a layer of locality-sensitive filtering, inspired by the quantum-walk algorithm for lattice sieving from [CL21]. We complement our asymptotic analysis of the quantum algorithms with numerical results, and observe that our quantum speed-ups for code sieving behave similarly as those observed in lattice sieving. In addition, we show that a natural quantum analog of the sieving-based ISD framework does not provide any speed-up over the first presented quantum ISD algorithm [Ber10]. Our analysis highlights that the framework should be adapted in order to outperform the state-of-the-art of quantum ISD algorithms [KT17, Kir18].
@@ 1.1 Main Contributions
We present several quantum algorithms for FindSolutions and analyze their asymptotic time and memory complexities. Our first quantum algorithm is a straightforward application of Grover’s algorithm, and serves as a baseline for comparison. Our other quantum algorithms are a quantum-walk algorithm and a variant thereof, where we apply the sparsification technique from [CL21]. Both quantum-walk algorithms make use of locality-sensitive filtering (LSF) to obtain a speed-up over Grover in the search for ‘solutions’. The use of LSF inside the quantum walk was first suggested by [CL21] in the context of lattice sieving, and one of our contributions is to show how to apply this layer of LSF in the context of codes (see Section 4.4). Specifically, we introduce a Hamming-metric variant of the ‘residual vectors’ used in [CL21], which behave somewhat differently than their Euclidean-metric analogs, but still allow us to apply LSF.
Besides providing an asymptotic analysis of the time and memory complexities of our quantum algorithms, we evaluate their performance numerically. Namely, the complexity of our quantum algorithms depends on some parameters that we optimize numerically to obtain the smallest runtime (and corresponding memory) for each of the algorithms. Our numerical results illustrate that we obtain a quantum speed-up over the classical algorithm from [DEEK24], and provide a comparison between our different quantum algorithms. In addition, we show that the quantum speed-ups we obtain align with the speed-ups observed in the state-of-the-art of lattice cryptanalysis (see Table 1). The Python code used for the numerical results of this work is available at https://github.com/lynnengelberts/quantum-sieving-for-codes-public.
@@ 2.4 Quantum Computation
We say that a quantum algorithm has time complexity $T$ if the circuit describing it uses at most $T$ elementary quantum gates. When $T$ is polynomial in the number of bits needed to write down the input, we say the algorithm is efficient. In the following, we define $[N]:=\{0,\ldots,N-1\}$.
@@
For the memory complexities of our algorithms, we will specify the sizes of the classical and quantum memories, and the sizes of the classical and quantum registers that the respective QRAM operations use. (That is, for an algorithm using the previous QRAM operations, we say it uses QRACM, respectively QRAQM, of size $N$.)
@@
For $N\in\mathbb{N}$, let $f\colon[N]\to\{0,1\}$ be a function and define $t:=|\{i\in[N]:f(i)=1\}|$. There exists a quantum algorithm, called Grover’s algorithm, that returns $i$ such that $f(i)=1$ with probability at least $2/3$, if such an $i$ exists, using $O(\sqrt{N/t})$ queries to $f$ and $\tilde{O}(\sqrt{N/t})$ elementary gates. The algorithm uses $\lceil\log_{2}(N)\rceil$ quantum memory.
@@
For $N\in\mathbb{N}$, let $f\colon[N]\to\{0,1\}$ be a function. Suppose there is an efficient quantum circuit $U_{f}$ that maps |$\ket{i}\to(-1)^{f(i)}\ket{i}$. Suppose $\mathcal{A}$ is an algorithm that can be implemented as a reversible quantum circuit and that returns $i\in[N]$ such that $f(i)=1$ with success probability $p>0$ in time $T$. Then there exists a quantum algorithm, called amplitude amplification, that returns $i\in[N]$ such that $f(i)=1$ with probability at least $\max(1-p,p)$ in time $O(1/\sqrt{p})\cdot T$.
@@ 3.1 Framework for Code Sieving
1 Sample a tower of codes $\{\mathbb{F}_{2}^{n}=\mathcal{C}_{0},\dots,\mathcal{C}_{n-k}=\mathcal{C}\}$, with dimension decrements of 1
2 Sample $N$ vectors independently and uniformly at random from $\mathcal{S}_{w}^{n}$ and add them to a list $\mathcal{L}_{0}$
@@
4       Invoke $\mathcal{A}_{\textsc{NNS}}$ on $\mathcal{L}_{i-1}$ to obtain $\mathcal{L}^{\prime}_{i}:=\{(\mathbf{x},\mathbf{y})\colon\mathbf{x},\mathbf{y}\in\mathcal{L}_{i-1},|\mathbf{x}+\mathbf{y}|=w\}$
@@
7                   Add $\mathbf{x}+\mathbf{y}$ to $\mathcal{L}_{i}$
@@
9      Discard some elements if $|\mathcal{L}_{i}|>N$
@@
By Lemma 3.1, if $N\geq{n\choose w}\big{/}\big{(}{w\choose w/2}{n-w\choose w/2}\big{)}$, then $\textsc{NNS}(n,w,N)$ has at least $N$ solutions in expectation. However, note that, in each iteration $i$ of Algorithm 1, on average only a quarter of the NNS solutions $(\mathbf{x},\mathbf{y})$ found by $\mathcal{A}_{\textsc{NNS}}$ will be added to the new list $\mathcal{L}_{i}$. Indeed, half of them are discarded as each pair is counted twice: $(\mathbf{x},\mathbf{y})$ and $(\mathbf{y},\mathbf{x})$. Secondly, the algorithm discards those $\mathbf{x}+\mathbf{y}$ that do not belong to $\mathcal{C}_{i}$. Since two subsequent codes $\mathcal{C}_{i-1},\mathcal{C}_{i}$ differ by one dimension, it follows that in expectation another half of the NNS solutions are discarded. Therefore, [DEEK24] propose to take $N\geq 4\cdot{n\choose w}\big{/}\big{(}{w\choose w/2}{n-w\choose w/2}\big{)}$ to maintain a list size of $N$ through all iterations in the sieving part.
@@ 3.2 NNS using Locality-Sensitive Filtering (LSF)
The underlying idea of locality-sensitive filtering (in $\mathbb{F}_{2}^{n}$) is the following: vectors $\mathbf{x},\mathbf{y}$ satisfying $|\mathbf{x}+\mathbf{y}|=w$, also known as near neighbors, can be found more efficiently if we restrict our search to local regions of $\mathbb{F}_{2}^{n}$. The algorithm proceeds as follows. It starts with covering the space $\mathbb{F}_{2}^{n}$ with (potentially overlapping) regions, each region corresponding to a certain center $\mathbf{c}\in\mathcal{C}_{f}$, for some $\mathcal{C}_{f}\subseteq\mathbb{F}_{2}^{n}$. Each vector $\mathbf{x}$ from the input list $\mathcal{L}$ is then filtered according to the centers, namely, it is inserted into a bucket corresponding to a center $\mathbf{c}$ if and only if $|\mathbf{x}\land\mathbf{c}|=\alpha$ for a fixed parameter $\alpha$ [^4] The analog in the lattice setting is that a unit vector is added to a bucket corresponding to some unit vector $\mathbf{c}\in\mathbb{R}^{n}$ if it has (absolute) inner product at least $\alpha$ with $\mathbf{c}$. This is what we refer to as the bucketing phase. Each bucket thus contains all the elements from $\mathcal{L}$ that belong to the same region. The algorithm then searches for near neighbors within each bucket, which we refer to as the checking phase. As the checking phase significantly affects the overall cost of the algorithm, our primary goal will be to improve its runtime by quantizing this part of the algorithm.
@@
The resulting NNS algorithm is given in Algorithm 2. It uses two subroutines: FindValidCenters and FindSolutions. For a given vector $\mathbf{x}\in\mathcal{L}$, the subroutine FindValidCenters returns the set $\mathcal{VC}_{\alpha}(\mathbf{x})=\{\mathbf{c}\in\mathcal{C}_{f}\ |\ |\mathbf{x}\land\mathbf{c}|=\alpha\}$. For a suitable choice of $\mathcal{C}_{f}$ (using so-called random product codes), the time complexity of this subroutine has been shown by [DEEK24] to be $|\mathcal{VC}_{\alpha}(\mathbf{x})|+2^{o(n)}$, as explained in the next section.
@@
3       $\mathcal{B}_{\alpha}(\mathbf{c})=\emptyset$
@@
7             Add $\mathbf{x}$ to $\mathcal{B}_{\alpha}(\mathbf{c})$
@@
9 $\mathcal{L}^{\prime}=\emptyset$
@@
11       Add $\textnormal{{FindSolutions}}(\mathcal{B}_{\alpha}(\mathbf{c}))$ to $\mathcal{L}^{\prime}$
@@
Depending on the size of $|\mathcal{C}_{f}|$, Algorithm 2 has to be repeated a certain number of times to find all solutions. The size of $|\mathcal{C}_{f}|$ also affects other factors of the time and memory complexity of Algorithm 2, and should thus be chosen carefully. The best time and memory complexity in [DEEK24, Cor. 4.2] are obtained by choosing $|\mathcal{C}_{f}|$ such that $\mathbb{E}[|\mathcal{VC}_{\alpha}(\mathbf{x})|]=2^{o(n)}$, ensuring that the expected runtime of FindValidCenters is $2^{o(n)}$. (An algorithm with similar time and memory complexities as [DEEK24, Cor. 4.2] was independently obtained by Carrier [Car20, Cor. 8.2.6].)
@@ 3.3 Random Product Codes and an Efficient Algorithm for FindValidCenters
The set of centers is then sampled uniformly at random from $R_{n,v,t,\kappa}$ to guarantee:
{outline}\1
[(1)] Efficient decodability: in some reasonable parameter regimes, one can compute $\mathcal{VC}_{\alpha}(\mathbf{x}):=\{\mathbf{c}\in\mathcal{C}\colon|\mathbf{x}\land\mathbf{c}|=\alpha\}$ in time asymptotically equal to its size $|\mathcal{VC}_{\alpha}(\mathbf{x})|$. \1[(2)] Random behavior: for $t$ not too large, a sample $\mathcal{C}$ behaves like a random code in the sense that the success probability that $\mathcal{C}\sim R_{n,v,t,\kappa}$ ‘captures’ a pair $(\mathbf{x},\mathbf{y})\in\mathcal{L}^{2}$ is the same (up to factors subexponential in $n$) as for a random code in $S_{v}^{n}$. [^5] This ‘randomness’ property puts a constraint on $t$. More precisely, as described in [DEEK24] (and in [BDGL16] for $\mathbb{R}^{n}$), we would like $t$ to be small enough to guarantee that we can approximate a cap/wedge in $\mathbb{F}_{2}^{n}$ by the Cartesian product of $t$ caps/wedges in $\mathbb{F}_{2}^{n/t}$. By Lemma 4.6 and Lemma 4.7 in [DEEK24], these approximations are satisfied up to a subexponential factor in $n$ if $t=o(n/\log(n))$.
@@ 4 Quantum Algorithms for NNS and Code Sieving
We recall the context in which FindSolutions is used. That is, consider the successful completion of the bucketing phase of Algorithm 2 for an input list $\mathcal{L}\subseteq\mathcal{S}_{w}^{n}$: for some bucketing parameter $\alpha$, we have sampled a set $\mathcal{C}_{f}\subseteq\mathcal{S}_{v}^{n}$ of bucket centers and created a data structure containing the buckets
$\mathcal{B}_{\alpha}(\mathbf{c}):=\{\mathbf{x}\in\mathcal{L}\colon|\mathbf{x}\land\mathbf{c}|=\alpha\}\subseteq\mathrm{Region}_{\alpha}(\mathbf{c}):=\{\mathbf{x}\in\mathcal{S}_{w}^{n}:|\mathbf{x}\wedge\mathbf{c}|=\alpha\}$ (1)
for all $\mathbf{c}\in\mathcal{C}_{f}$. Then the goal of FindSolutions is to find all $\mathbf{x},\mathbf{y}\in\mathcal{B}_{\alpha}(\mathbf{c})$ satisfying $|\mathbf{x}+\mathbf{y}|=w$.
@@
This problem is a variant of the near-neighbor search problem (Problem 3.2) defined in Section 3. An important difference here is that the input list $\mathcal{B}_{\alpha}(\mathbf{c})$ is not uniformly random on $\mathcal{S}_{w}^{n}$ but on the set $\{\mathbf{x}\in\mathcal{S}_{w}^{n}:|\mathbf{x}\wedge\mathbf{c}|=\alpha\}$. Consequently, the probability (denoted $p$ below) that a uniformly random pair from the input list forms a solution pair does not only depend on $n,w$, but also on the parameters $v,\alpha$, as further shown in the next section.